Nicole:

Which is not a step when constructing a line parallel to the x-axis of a coordinate plane through a point? Measure the distance from the point to the x-axis Place the compass on the point and mark an arc on both sides of the x-axis. Connect the arc intersections using a straight edge. Draw a transversal through the x-axis and a point not on the x-axis.

Nicole:

@dude

dude:

Ideas? https://www.mathopenref.com/constparallel.html Can help if you're unsure of the steps

Nicole:

im thinking its d?

dude:

No no They drew the transversal over the point P and the x axis, so that is a step in the construction of a life (Note on how the ruler was used measurement vs straightedge)

Nicole:

A?

Nicole:

or C

dude:

*It would be A* Because they didn't measure anything They only used the ruler to draw straight lines (it was used only as a straightedge)

dude:

(C was the last step in the construction)

Nicole:

Got it okay

Nicole:

If the midpoint between (18, y) and (20, -15) is (19, -5), find the value of y.

dude:

Make a new post after this one

Nicole:

okay

dude:

This is harder said than done, so I'll set it up for you, if you dont understand something, feel free to ask Use the midpoint formula \((\dfrac{x_1+x_2}{2}, \dfrac{y_1+y_2}{2})=(x_m,y_m)\) I am using m as a letter to define "middle" -- this is informal (18, y) -> (\(x_1,y_1\)) (20,-15) -> (\(x_2,y_2\)) (19,-5) -> (\(x_m,y_m\)) Substitute \((\dfrac{18+20}{2}, \dfrac{y-15}{2})=(19,-5)\) It could help to split it into x and y \(\dfrac{18+20}{2}=19\) \(\boxed{\dfrac{y-15}{2}=-5}\) We only need to solve for y Do you know how to solve for y from here?

Nicole:

y=5

dude:

Ye good

dude:

You can check this if you want Do you want me to show you how to check? Or are you okay?

Nicole:

no thank you I got that one. Im going to make a new post